OpenStudy (anonymous):

HELP!! Write a rule for the translation of point A to point A' and use that rule to locate points B' and C' A(-1,-2) A'(-4 0) B(1 2) C(3 -1)

OpenStudy (anonymous):

I think its x+2 y-4

OpenStudy (anonymous):

for translation, the points all move the same distance and the same direction. First we determine how much distance did these points move by. so A(-1,-2) moves to A'(-4 0) therefore wat is the change in the x and y value of point A?

OpenStudy (anonymous):

x+2 y-4

OpenStudy (anonymous):

no, lets look at the x value for A and A' which are -1 and -4 respectfully.

OpenStudy (anonymous):

jki fund it lol

OpenStudy (anonymous):

oh so isit x-4 y+2

OpenStudy (anonymous):

cool :) now wat was the change in A's x value?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ill stop jumpin ahead

OpenStudy (anonymous):

ok wat do u do to -1 to get -4?

OpenStudy (anonymous):

move over4 from A

OpenStudy (anonymous):

that would be so if A's x value was 0 cause 0 - 4 = -4

OpenStudy (anonymous):

oh so u move to the left 4 and up2

OpenStudy (anonymous):

ur close, Pt A moves up 2 and to the left but doesnt move to the left by 4 but by 3

OpenStudy (anonymous):

oh I just counted wrong lol so then would it be x-3 y+2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and then use that to get B' and C'

OpenStudy (anonymous):

yes, so do that and ill let u know if ur right :)

OpenStudy (anonymous):

kk I got B'=(-2 5) C'(0 2)....oh ya I have a cople more questions for this one if you wouldn't mind helping on them when I get this one done that would be cool of u :)

OpenStudy (anonymous):

no,forpt B(1 2) what you do is minus -3 from B's x value 1, and add 2 to B's y value 2. do over C as well, and sure id love to help :)

OpenStudy (anonymous):

coo ok thanks

OpenStudy (anonymous):

B'=(-2 4) C'(0 1)

OpenStudy (anonymous):

wait my bad it's right

OpenStudy (anonymous):

so next question :)

OpenStudy (anonymous):

ok I was gonna say lol then I have to wite a rule to translate A'B'C' back to ABC but would that just be x+3 y-2

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

nex question:

OpenStudy (anonymous):

D'(5 -2) E'(7 -1) the image of (xy) (x+1)(y-5) find DE

OpenStudy (anonymous):

and write the vector that translates DE to D'E'

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

kkk

OpenStudy (anonymous):

back

OpenStudy (anonymous):

cool u got the questions I typed

OpenStudy (anonymous):

wat's (xy) (x+1)(y-5)?

OpenStudy (anonymous):

oh I figured them ones out lol

OpenStudy (anonymous):

i still dont understand :s

OpenStudy (anonymous):

it gives you how to find D'E' u just substitute add wit subtraction and get DC so its (x-1)(y+5) instead of (x+1)(y-5)

OpenStudy (anonymous):

A golfer teein off on an easy par 3 hole from point A hits the ball to point B the second stroke moves the ball from point B to C(on the green) ten feet away from the hole(point D(4 3)/. Write the component for AB BC and AD

OpenStudy (anonymous):

do I just use the distance formula

OpenStudy (anonymous):

u have the coordinates of A, B and C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think so.. im not sure wat the word component means

OpenStudy (anonymous):

if by component they mean the length then yes

OpenStudy (anonymous):

yea I think so give me asec to find and solve and ill let u kno what I get

OpenStudy (anonymous):

wat's the formula again? lol i forgot :P

OpenStudy (anonymous):

AB=(2-x1)^2+(y2-y1)^2 all over sqrt

OpenStudy (anonymous):

\[AB=\sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (anonymous):

ok cool :) now do it and ill tell u if ur right :)

OpenStudy (anonymous):

ok for AB=7.21 Bc+3.31 Ad+8.54

OpenStudy (anonymous):

u have to give me the coordes for A, B and C for me to check ur work

OpenStudy (anonymous):

theyr at the verytop in my first question

OpenStudy (anonymous):

a(-1 -2)b(1 2)c(3 -1) d(4 3)

OpenStudy (anonymous):

say it again cause i dont see it..

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

crap, brb

OpenStudy (anonymous):

kkk

OpenStudy (anonymous):

back

OpenStudy (anonymous):

doing it now

OpenStudy (anonymous):

cool cool di du get the answers

OpenStudy (anonymous):

AB = 4.47

OpenStudy (anonymous):

\[\sqrt{(1-(-1))^{2}+(2-(-2))^{2}}\]

OpenStudy (anonymous):

show me ur working for AB

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

\[\sqrt{(-2-4)^{2}+(2-0)^{2}}\] \[\sqrt{36+15}\] \[\sqrt{52}or7.21\]

OpenStudy (anonymous):

I did it wrong

OpenStudy (anonymous):

:) correct it and gud job!

OpenStudy (anonymous):

gud job for seeing that something is wrong

OpenStudy (anonymous):

ok so now I got AB=4.47 BC=3.61 and AD=7.07

OpenStudy (anonymous):

i only did AB and ur answer is right :) ill now do BC and AD

OpenStudy (anonymous):

I did I gave u my answers above^^^

OpenStudy (anonymous):

BC is correct as well :)

OpenStudy (anonymous):

this is like the only thing im good at lol

OpenStudy (anonymous):

AD is correct :) so all ur answer r correct :D

OpenStudy (anonymous):

ihave 3 mor questins lol next question: if the golfer makes a ten foot putt from from point C into the hole (point D)the how far did the ball trave en route from A to B to C to D.(hint:use the paythagoren troples 3,4,5 5,12,13,and 8,15,17 and a scale of 1 unit=10feet)

OpenStudy (anonymous):

jk only 2 :)

OpenStudy (anonymous):

haha cool :P

OpenStudy (anonymous):

... i cannt picture the triangle here. wat i see is this